A revolver of mass 5kg releases a bullet of mass 0.10kg at a speed of 100m/s.find the revolvers speed recoil
1
Expert's answer
2020-01-20T05:18:34-0500

Due to Newton’s Third Law, every action has an equal and opposite reaction. The statement means that in every interaction, there is a pair of forces acting on the two interacting objects. The size of the forces on the first object equals the size of the force on the second object.

So, 0.1kg of bullet with speed 100 m/s and 5kg gun recoil with speed "\nu".

"0.1 * 100 = 5 * \nu;"

"\nu = (0.1 * 100) /5 = 10/5 = 2 m/s"

Answer: "2 m\/s"